# Passing the Buck to AI: How Individuals’ Decision-Making Patterns Affect Reliance on AI

### v1

@@ 3.2. Experiment Design
To examine our hypotheses, we designed an online study (preregistered at link) that includes (i) a questionnaire and (ii) a task requiring participants to decide on the factuality of a series of nutrition statements while having the option to seek out AI suggestions (AI decision and AI explanations). We generated AI decisions and explanations using ChatGPT before the study (rather than having participants interact with ChatGPT live) to ensure consistency of response quality. The study was approved by our Institutional Review Board and launched in English on the volunteer-based online study platform [AnonymousOnlineLab]222name changed to preserve anonymity. We encouraged honest participation by informing participants in advance that they would receive personalized result—their own decision-making patterns and performance on evaluation food facts and myths—at the end of the study.
Nutrition Statements for the Main Task: We used 30 nutrition statements and their accuracy (i.e., fact or myth) from Florença et al. (2021) who had curated a list of popular food (mis)-conceptions from online sources. In their study with 503 participants, several of these statements were inaccurately classified as facts or myths by participants, and even by participants working in areas related to nutrition, suggesting that the statements have a range of difficulty. Among the 30 statements, 9 nutrition statements are facts and 21 are myths. One example false statement (a myth) is “Pregnant women should be eating for two”, while an example true statement (a fact) is “Dairy products should be consumed in between two and three portions per day”. The full list of statements with fact/myth labels is in the Supplementary Materials.
@@ 3.3. Study Procedure
If participants chose “I’d like to see the AI decision,” they were provided two additional options: to see an AI explanation or to immediately accept the AI decision (without seeing an explanation). This design was motivated by our hypotheses: see AI decisions is designed to observe whether users prefer to refer to the AI’s decision before making their own. Subsequently, we offer the option to see AI explanations to investigate whether some users may want to further investigate and seek additional details.
After an AI decision is revealed, participants could select Use the AI decision to indicate they accept the AI decision; however, this design revealed its limitations as participants barely selected this option based on our analysis of results which we discuss later. For each statement, we recorded the time they took to evaluate the AI’s decision and explanations before final decision-making.
@@ 3.4. Metrics
  • Time Spent Seeing AI Explanations is a continuous variable that measures the time it takes participants to select a final evaluation of a statement after they choose to see the AI explanation. We will only examine this variable for participants who select a response after they click “See AI Explanation”.
  • Perceived AI Reliance measures individuals’ self-reported reliance on AI assistance in the study based on their responses to the question “How much did you rely on the AI in your decisions for food facts and myths?” We coded levels 1 to 5 (1 for “Didn’t rely on it at all”, 5 for “I relied on it all the time”). We adopted a one-item measurement approach to examine individuals’ perception-based reliance on AI based on prior study design (Cao and Huang, 2022). The item was adapted to focus specifically on our nutrition task, assessing participants’ reliance explicitly during the study.
We examined the effects of decision-making patterns while considering the demographic covariates (age, gender, education levels) as well as the following variables:
  • Vigilance, Hypervigilance, Buckpassing: Based on participants’ answers to the Melbourne Decision Making Questionnaire (MDMQ) (Mann et al., 1997) on a 3-point scale (“True for me” (score 2), “Sometimes true for me” (score 1), “Not true for me” (score 0)), we obtained the total score for each of their decision-making dimensions. Each decision-making pattern has its own independent subscale, consisting of a unique set of question statements distinct from those used in the other subscales. Therefore, while we excluded procrastination in our hypotheses and did not administer that portion of the MDMQ, this should not impact the validity of measuring buckpassing, vigilance, and hypervigilance. The maximum score for hypervigilance is 10, and for vigilance and buckpassing, it is 12. The vigilance and buckpassing subscales consist of 6 items each, while the hypervigilance subscale consists of 5 items.
@@ 3.5. Participants
Participants in our study were 917 volunteers recruited via the online study platform [AnonymousOnlineLab] [^4] Name omitted for anonymity. between August 2023 and April 2024. The platform was chosen to allow us to recruit diverse participants of various ages and education levels to ensure a broad spread of decision-making styles. Participation on [AnonymousOnlineLab] is open to anyone without signing up; we therefore obtained a waiver of parental consent from our IRB so that minors were able to take our study. To reduce the risk of participants mistaking factually incorrect nutrition information as correct, the study instructions clearly stated that not all nutrition statements are factual and that the goal is to test how much participants know about nutrition. As part of this debrief, participants were also told which of the nutrition statements they wrongfully assumed to be true at the end of the experiment.
@@
The final set of participants reported being from 69 countries, with the majority from the United States (44%), followed by the United Kingdom (9%), Canada (8%), India (5%), Germany (3%), and Australia (3%, all others $\leq 3\%$). Half of our participants reported pursuing or having obtained a college education, approximately 29% of participants reported pursuing or obtaining a high school education; with the rest reporting to pursue or having obtained a graduate education. Our study participants were between 12 to 90 years of age ($M=31$, $SD=14$). A detailed breakdown of demographics for participants is shown in Table  3.
@@ 4.1. Descriptive Statistics
The full summary statistics of decision-making patterns for the different demographics are shown in the Appendix (Table 9). On average, our study participants scored 9.41 ($SD=2.31$) out of 12 on vigilance, 4.25 ($SD=2.61$) out of 10 on hypervigilance, and 4.56 ($SD=3.03$) out of 12 on buckpassing. Each subscale reaches a satisfactory level of internal reliability. Cronbach’s alphas for each subscale is 0.74 (vigilance), 0.81 (buckpassing), and 0.74 (hypervigilance).  [^5] Cronbach’s alpha measures internal consistency and reliability of a scale, ranging from 0 to 1. Values from 0.70-0.90 are recommended to be acceptable values (Tavakol and Dennick, 2011). (Cronbach alphas were 0.80,0.87,0.74 from Mann et al. (1998)). The mean and standard deviation of our measures fall within a similar range compared to those observed in Mann et al. (1998) across different countries. Participants’ hypervigilance has significantly positive correlation with their measure of buckpassing ($r=0.59,p<0.001$), which aligns with prior studies (Filipe et al., 2020; Mann et al., 1997). Additional factor analysis confirms that these three different decision-making patterns are identified with the same corresponding question items from Mann et al. (1998). Details regarding factor loadings and a corresponding diagram are included in the Appendix. These similarities suggest the reliability of our measure of decision-making patterns and the robustness of our samples.
@@
On average, participants accurately evaluated 70% of the statements ($M=70\%,SD=14\%$). The performance of our participants is significantly correlated with participants in Florença et al. (2021) ($r=0.58,p<0.001$), indicating that most of our participants exhibit a similar knowledge pattern regarding common food myths and facts.
Performance did not significantly differ across the different decision-making styles.
@@
4.2. Effect of Decision-Making Patterns on Seeing AI Decisions
@@ 4.3. Effect of Decision-Making Patterns on Seeing AI Explanations
Next, we examined the effect of decision-making patterns on how much time people spend reviewing AI explanations, if they choose to see them. On average, participants spent 14 seconds seeing AI explanations ($M=14.34,SD=10.73$). We used a linear mixed-effects regression model, considering that each participant could have their own reading speed and each statement has different lengths of AI explanations. Per Table 5, buckpassing has a statistically significant negative effect on individuals’ time spent evaluating AI explanations ($\beta=-0.40,p<0.05$). For one additional score of increase in buckpassing dimension, the time users spend on seeing explanations decreases by 0.4 seconds. This suggests individuals who score higher in buckpassing tend to spend less time evaluating AI explanations before they make their final decisions. On the contrary, vigilance has a statistically significant positive effect on participants’ time spent evaluating AI explanations ($\beta=0.53,p<0.05$). For one score of increase in vigilance dimension, the time users spend on reading explanations increases by 0.53 seconds. These results support our hypotheses H1b and H3b, but not H2b.
@@
Domain_Knowledge (Average) --0.162 (1.035)
@@
Pursuing or Have obtained graduate education --0.066 (1.431)
@@
Note: p$<$0.1; ∗∗p$<$0.05; ∗∗∗p$<$0.01
@@ 4.4. Effects of Decision-making Patterns on Perceived AI Reliance
To test H1c, H2c, and H3c, we constructed an ordinal logistic regression model while controlling for participants’ demographic covariates including age, education level, domain knowledge, and perception of AI. Ordinal logistic regression is often used to predict dependent variables that can be ordered in a natural way such as mild, moderate, severe (Harrell, 2015). Per Table 6, our result indicates the effect of buckpassing on perceived AI reliance is statistically significant ($OR=1.09,95\%\text{ CI}[1.03,1.15],p<0.05$). For one score of increase in buckpassing, the odds of having more self-perceived reliance on the AI increases by 9%, suggesting that people who score high in buckpassing are more likely to rely on AI when making their decisions. Figure 3 shows the distribution of response in terms of level of reliance related to individuals’ buckpassing tendency. Noteworthily, 12% of people who received a low buckpassing score reported to rely on the AI “all the time” (rating of 5) or most of the time (rating of 4), while 20% of those who received a high buckpassing score reported the same.
No effect of vigilance or hypervigilance is observed. Hypotheses H1c and H2c are therefore not supported, while H3c is supported.
Figure 3. Perceived AI Reliance Variations Across Different Levels of Buckpassing: individuals who score high in buckpassing report a higher reliance on AI.
@@ 4.5. Who tends to score high in buckpassing?
Figure 4. Examining the demographic distribution of buckpassing tendency: participants who have a high school education or below score higher in buckpassing (23% in high buckpassing level); participants who are aged under 18 and in their early 20s also tend to score higher in buckpassing.
Our analysis indicates that a buckpassing tendency significantly influences participants’ use of AI suggestions and reliance on AI. Thus we further explore the demographic distribution of buckpassing scores among our participants to gain insights into which demographic groups exhibit a higher buckpassing tendency. Figure 4 shows the distribution of buckpassing levels among different education levels and age groups. Age is treated as a categorical variable in this descriptive analysis to obtain a more granular understanding of age group differences. Teenagers and young adults in our sample tend to score higher in buckpassing compared to participants who are aged above 35. Participants who obtain a lower level of education (high school education or below) also exhibit a higher buckpassing tendency compared to those with a higher level of education (pursuing or having obtained college or graduate education).
@@ 5. Discussion
Individuals’ decision-making patterns have been shown to influence their decision outcomes and quality of life (Kim et al., 2022; Deniz, 2006), underlining the importance of understanding how people vary in making decisions. However, despite the increasing role of AI in decision-making contexts, how individual cognitive decision-making processes influence the use of AI technology has been unknown. Our study offers evidence of this influence; we discuss our findings and consider their implications for designing AI systems and explainable AI research below.
@@
In this work, we extend the current understanding of factors behind individuals’ reliance on AI. Consistent with our hypotheses, we found that people who tend to defer decision-making patterns to others (score high in buckpassing) are more likely to seek out suggestions from an AI and report relying on it, though they spend less time reading AI’s explanations. While the difference can be described as a small to medium effect—those who score low on buckpassing viewed the AI’s decisions 25% of the time versus 33% for those who scored high on buckpassing—this difference can be profound when AI decision-making tools are frequently used. The difference may also gain importance with the number of decisions to be made as decision fatigue sets in (Chen et al., 2023) or as task difficulty or uncertainty increases (Vasconcelos et al., 2023; Salimzadeh et al., 2024), leading to a higher likelihood of overreliance.
@@
Prior work has identified individual-related factors that affect individuals’ trust and reliance on AI including personal traits and demographic characteristics. In addition to personality and self-confidence (Cai et al., 2022; Chong et al., 2022), our work suggests that individuals’ decision-making patterns could affect individuals’ interactions with AI systems. We encourage future work to examine these factors in various kinds of human-AI interactions. In addition, prior HCI research shows that individuals’ trust and reliance vary by their age, education levels, and perception of AI (Ma and Huo, 2023; Araujo et al., 2020). Our work extends this by showing that individuals’ buckpassing tendency has a marginal effect on AI use when accounting for these demographic covariates. While the significance of the effect decreases after controlling these factors, it is probably due to the correlation between these demographic factors and decision-making patterns (see correlation matrix table in the Appendix).
@@
What might explain these findings is the stress level and confidence people experience while making decisions. Mann et al. (1997)’s work has indicated that these decision-making tendencies can be seen as different kinds of coping mechanisms during decision conflict. Individuals with a strong buckpassing tendency often experience high psychological stress during decision-making tasks, while vigilant decision-makers experience a lower level of stress (Mann et al., 1997). Experiencing high psychological stress while making decisions might propel people to seek out an immediate relief by turning to AI suggestions. Mann et al. (1998) also identified that individuals who score high in buckpassing tend to have a lower confidence in their own decision-making. This could potentially propel participants in our study to seek AI’s suggestions when they were not confident about making the decisions themselves. Interestingly, we did not observe any effect of vigilance or hypervigilance on participants’ tendency to see AI suggestions. This might be due to the low variance of participants on vigilance ($SD=2.31$) and hypervigilance ($SD=2.61$) compared to buckpassing pattern ($3.03$). However, we observed participants who score high in vigilance spent significantly more time seeing AI explanations. This suggests that an increased vigilance goes hand in hand with careful reading of any explanations people are provided, likely as a tool for evaluating whether the AI’s decision can be reasonably trusted.
@@
AI-assisted decision-making could yield both positive and negative outcomes for people with a tendency for buckpassing. On the one hand, the rise of AI systems could ease the burden of information seeking and processing for individuals to make effective decisions, if the performance of these systems is reliable. As evidenced by prior work, making decisions with the help of AI could lead to more accurate decisions when the AI has high advice accuracy and quality (Vodrahalli et al., 2021; Tschandl et al., 2020) (though there are some exceptions where this was found not to be the case (Vaccaro and Waldo, 2019; Jacobs et al., 2021)). AI assistance can also lower mental demand (Buçinca et al., 2021). With highly reliable AI systems, AI suggestions could help relieve the decision-making stress experienced by people with a buckpassing tendency. On the other hand, given the current risks of AI hallucinations, where AI generates unfaithful or nonfactual content (Bubeck et al., 2023), and AI overconfidence (Zhou et al., 2024), where AI outputs sound confident yet are inaccurate (Zhou et al., 2024), buckpassers are more susceptible to overreliance on AI, potentially resulting in misleading decisions and negative decision outcomes.
The benefits and risks of AI-assisted decision-making might be exacerbated for some people and groups, as our study and other research show that decision-making patterns vary across demographics. Consistent with prior findings (Filipe et al., 2020), our study suggests that buckpassing tendency is more prevalent among younger people (especially individuals under 18) and people with high school education or below. This warrants greater intervention to guide these populations in their usage of AI systems. Prior work from Mann et al. (1998) indicates that people from East Asian countries (e.g., Japan, Hong Kong) exhibit a higher buckpassing tendency compared to people in Western countries (e.g., USA, Australia, New Zealand). East Asians have also been found to have a lower decision-making self-esteem, which is generally in line with a preference for making decisions in a group (Mann et al., 1998). These demographic groups may seek out information from the AI more often, benefiting more from AI systems that have high accuracy but also facing greater risks of overreliance.

### v2

@@ 3.2. Experiment Design
To examine our hypotheses, we designed an online study (preregistered at link) that includes (i) a questionnaire and (ii) a task requiring participants to decide on the factuality of a series of nutrition statements while having the option to seek out AI suggestions (AI decisions and AI explanations). We generated AI decisions and explanations using ChatGPT before the study (rather than having participants interact with ChatGPT live) to ensure consistency of response quality. The study was approved by our Institutional Review Board and launched in English on the volunteer-based online study platform LabintheWild [^3] https://labinthewild.org/. We encouraged honest participation by informing participants in advance that they would receive personalized result—their own decision-making patterns and performance on evaluation food facts and myths—at the end of the study.
Nutrition Statements for the Main Task: We used 30 nutrition statements and their accuracy (i.e., fact or myth) from Florença et al. (2021) who curated a list of popular food (mis)-conceptions from online sources. In their study with 503 participants, several of these statements were inaccurately classified as facts or myths by participants, and even by participants working in areas related to nutrition, suggesting that the statements have a range of difficulty. Among the 30 statements, 9 nutrition statements are facts and 21 are myths. One example false statement (a myth) is “Pregnant women should be eating for two”, while an example true statement (a fact) is “Dairy products should be consumed in between two and three portions per day”. The full list of statements with fact/myth labels is in the Supplementary Materials.
@@ 3.3. Study Procedure
This figure illustrates the interaction flow of the main task for participants. The figure indicates two possible interaction flows for participants when they are evaluating food facts or myths. One of the interaction flows is that participant can choose to provide their answer based on their own knowlege and select the correponding answer. The other workflow is particpants can choose to select See AI decision button to see AI’s decision for the evaluation of statement. After the AI’s decision is revealed, participants would be shown a button that says ”Show me the AI’s explanation”. At this point, they can either skip the explanation and just make their final evaluation, or click the button to reveal AI’s text-based explanations.
@@
If participants chose “I’d like to see the AI decision,” they were provided two additional options: to see an AI explanation or to immediately accept the AI decision (without seeing an explanation). This design was motivated by our hypotheses: see AI decision is designed to observe whether users prefer to refer to the AI’s decision before making their own. Subsequently, we offered the option to see AI explanations to investigate whether some users might want to further investigate and seek additional details.
After an AI decision was revealed, participants could select Use the AI decision to indicate they accept the AI decision; however, this design revealed its limitations as participants barely selected this option based on our analysis of results which we discuss later. For each statement, we recorded the time they took to evaluate the AI’s decision and explanations before final decision-making.
@@ 3.4. Metrics
  • Time Spent Seeing AI Explanations is a continuous variable that measures the time it takes participants to select a final evaluation of a statement after they choose to see the AI explanation. We only examined this variable for participants who select a response after they click “See AI Explanation”.
  • Perceived AI Reliance measures individuals’ self-reported reliance on AI assistance in the study based on their responses to the question “How much did you rely on the AI in your decisions for food facts and myths?” We coded levels 1 to 5 (1 for “Didn’t rely on it at all”, 5 for “I relied on it all the time”). We adopted a one-item measurement approach to examine individuals’ perception-based reliance on AI based on prior study design (Cao and Huang, 2022). The item was adapted to focus on our nutrition task, assessing participants’ reliance explicitly during the study.
We examine the effects of decision-making patterns while considering the demographic covariates (age, gender, education levels) as well as the following variables:
  • Vigilance, Hypervigilance, Buckpassing: Based on participants’ answers to the Melbourne Decision Making Questionnaire (MDMQ) (Mann et al., 1997) on a 3-point scale (“True for me” (score 2), “Sometimes true for me” (score 1), “Not true for me” (score 0)), we obtained the total score for each of their decision-making dimensions. Each decision-making pattern has its own subscale, consisting of a unique set of statements distinct from those used in the other subscales. Therefore, while we excluded procrastination in our hypotheses and did not administer that portion of the MDMQ, this should not impact the validity of measuring buckpassing, vigilance, and hypervigilance. The maximum score is 10 for hypervigilance, and 12 for both vigilance and buckpassing. The vigilance and buckpassing subscales consist of 6 items each, while the hypervigilance subscale consists of 5 items.
@@ 3.5. Participants
Participants in our study were 917 volunteers recruited via the online study platform LabintheWild between August 2023 and April 2024. The platform was chosen to allow us to recruit diverse participants of various ages and education levels to ensure a broad spread of decision-making styles. Participation on LabintheWild is open to anyone without signing up; we therefore obtained a waiver of parental consent from our IRB so that minors were able to take our study. To reduce the risk of participants mistaking factually incorrect nutrition information as correct, the study instructions clearly stated that not all nutrition statements are factual and that the goal is to test how much participants know about nutrition. As part of this debrief, participants were also told which of the nutrition statements they wrongfully assumed to be true at the end of the experiment.
@@
The final set of participants reported being from 69 countries, with the majority from the United States (44%), followed by the United Kingdom (9%), Canada (8%), India (5%), Germany (3%), and Australia (3%, all others $\leq 3\%$). Half of our participants reported pursuing or having obtained a college education, approximately 29% of participants reported pursuing or obtaining a high school education; with the rest reporting to pursue or having obtained a graduate education. Our study participants were between 12 to 90 years of age ($M=31$, $SD=14$). A detailed breakdown of demographics for participants is shown in Table 3.
@@
This table shows the sample sizes of study participants across different demographics, including their education levels, gender, age group, self-reported knowledge of nutrition, and their perception of the accuracy of AI-generated information. The table contains columns for the different categories and the corresponding sample sizes for each.
@@ 4.1. Descriptive Statistics
The full summary statistics of decision-making patterns for the different demographics are shown in the Appendix (Table 9). On average, our study participants scored 9.41 ($SD=2.31$) out of 12 on vigilance, 4.25 ($SD=2.61$) out of 10 on hypervigilance, and 4.56 ($SD=3.03$) out of 12 on buckpassing. Each subscale reaches a satisfactory level of internal reliability. Cronbach’s alphas for each subscale is 0.74 (vigilance), 0.81 (buckpassing), and 0.74 (hypervigilance).  [^5] Cronbach’s alpha measures internal consistency and reliability of a scale, ranging from 0 to 1. Values from 0.70-0.90 are recommended to be acceptable values (Tavakol and Dennick, 2011). Cronbach’s alphas were 0.80,0.87,0.74 from Mann et al. (1998). The mean and standard deviation of our measures fell within a similar range compared to those observed in Mann et al. (1998) across different countries. Participants’ hypervigilance had significantly positive correlation with their measure of buckpassing ($r=0.59,p<0.001$), which aligns with prior studies (Filipe et al., 2020; Mann et al., 1997). Additional factor analysis confirmed that these three different decision-making patterns were identified with the same corresponding question items from Mann et al. (1998). Details regarding factor loadings and a corresponding diagram are included in the Appendix. These similarities suggest the reliability of our measure of decision-making patterns and the robustness of our samples.
@@
On average, participants accurately evaluated 70% of the statements ($M=70\%,SD=14\%$). Participants who scored high in buckpassing (M = $68\%$) had a slightly lower accuracy than those who scored average (M = $70\%$) and low in buckpassing (M = $71\%$). When AI decisions were incorrect, participants scored much lower in accuracy ($M=23\%$) when they chose to see AI decisions compared to times when they did not ($M=57\%$). When AI decisions were correct, participants scored slightly higher in accuracy when they chose to see AI decisions ($M=90\%$) compared to times when they didn’t ($M=84\%$). A detailed breakdown of accuracy across different conditions and buckpassing level is included in the appendix (Figure 10). Performance did not significantly differ across the different decision-making styles. The performance of our participants on each statement was significantly correlated with participants in Florença et al. (2021) ($r=0.58,p<0.001$), indicating that most of our participants exhibit a similar knowledge pattern regarding common food myths and facts.
Agreement with AI suggestions
To examine participants’ decision behaviors after they see AI decisions, we measured their agreement rate. Proportion of agreement with AI suggestions is often used as a behavioral metric of reliance in prior studies where AI suggestions are automatically shown (Lu et al., 2024; Cao and Huang, 2022; Cao et al., 2023). However, since our study design provided AI suggestions on demand, we measured participants’ conditional agreement rate conditioned on whether they chose to see AI decisions. Participants agreed with AI decisions on about 83% of statements when they chose to see them, compared to 67% when they did not. Among all participants, around 49% (395 out of 810) of participants agreed with AI every time they chose to see AI decisions. We also visualized the distribution of participants’ individual conditional agreement rates in the Appendix (Figure 9).
@@
This figure is a bar chart displaying the relationship between buckpassing (a decision-making style) and the frequency of users choosing to ”See AI Decisions” across three buckpassing groups: Low, Average, and High. On the y-axis is the average frequency of selecting ”See AI Decisions” and the error bar. The scale ranges from 0.0 to 0.4. On the x-axis is the buckpassing levels divided into three categories: low, average, and high. Low level includes participants whose buckpassing score is 1 standard deviation below the mean (marked in red). Average level includes participants with an average buckpassing score. High level includes participants whose buckpassing score is 1 standard deviation above the mean. Sample size for each group are displayed above the bars: Low buckpassing group has n = 139. Average buckpassing group has n = 540. High buckpassing group has n = 132. The heights of the bars show that participants in the Low and High buckpassing groups selected ”See AI Decisions” less frequently than those in the Average group. The error bars are 95% confidence intervals for the mean frequencies. The Low buckpassing group has a slightly lower frequency with a narrow confidence interval, while the Average and High groups have higher frequencies with overlapping confidence intervals.
4.2. Effect of Decision-Making Patterns on Seeing AI Decision
@@ 4.3. Effect of Decision-Making Patterns on Seeing AI Explanations
Next, we examined the effect of decision-making patterns on how much time people spend reviewing AI explanations, if they choose to see them. On average, participants spent 14 seconds seeing AI explanations ($M=14.34,SD=10.73$). We used a linear mixed-effects regression model, considering that each participant could have their own reading speed and each statement had different lengths of AI explanations. Per Table 5, buckpassing had a statistically significant negative effect on individuals’ time spent evaluating AI explanations ($\beta=-0.40,p<0.05$). For one additional score of increase in buckpassing dimension, the time users spend on seeing AI explanations decreased by 0.4 seconds. This suggests individuals who score higher in buckpassing tend to spend less time evaluating AI explanations before they make their final decisions. On the contrary, vigilance had a statistically significant positive effect on participants’ time spent evaluating AI explanations ($\beta=0.53,p<0.05$). For one score of increase in vigilance dimension, the time users spent on reading explanations increased by 0.53 seconds. These results support our hypotheses H1b and H3b, but not H2b.
@@
Domain_Knowledge (Average) -0.162 (1.035)
@@
Pursuing or Have obtained graduate education -0.066 (1.431)
@@
Note: p<<0.1; ∗∗p<<0.05; ∗∗∗p<<0.01
This image shows the results of a linear mixed model predicting individuals’ time spent evaluating AI explanations. The table displays the coefficient (standard error) for each independent variable.The independent variables include Buckpassing, Vigilance, Hypervigilance, Age, Domain_Knowledge (Average), Domain_Knowledge (Below Average), Pursuing or Have obtained college education, Pursuing or Have obtained graduate education, Perception of AI (High), and Constant. The table also provides the number of observations and the Marginal R2 / Conditional R2 values. The results indicate that participants’ buckpassing and vigilance tendency significantly affect the time they spent reading AI’s explanations if they chose to see them. Specifically, higher buckpassing is associated with less time spent evaluating AI explanations, while higher vigilance is associated with more time spent evaluating AI explanations
@@ 4.4. Effects of Decision-making Patterns on Perceived AI Reliance
To test H1c, H2c, and H3c, we constructed an ordinal logistic regression model while controlling for participants’ demographic covariates including age, education level, domain knowledge, and perception of AI. Ordinal logistic regression is often used to predict dependent variables that can be ordered in a natural way such as mild, moderate, severe (Harrell, 2015). Per Table 6, our result indicates the effect of buckpassing on perceived AI reliance was statistically significant ($OR=1.09,95\%\text{ CI}[1.03,1.15],p<0.05$). For one score of increase in buckpassing, the odds of having more self-perceived reliance on the AI increases by 9%, suggesting that people who score high in buckpassing are more likely to rely on AI when making their decisions. Figure 3 shows the distribution of response in terms of level of reliance related to individuals’ buckpassing tendency. Noteworthily, 12% of people who received a low buckpassing score reported to rely on the AI “all the time” (rating of 5) or most of the time (rating of 4), while 20% of those who received a high buckpassing score reported the same.
No effect of vigilance or hypervigilance was observed. Hypotheses H1c and H2c are therefore not supported, while H3c is supported.
Figure 3. Perceived AI Reliance Variations Across Different Levels of Buckpassing: individuals who scored high (one standard deviation above average among participants) in buckpassing reported a higher reliance on AI compared to those who scored low (one standard deviation below average) in buckpassing.
[…]
@@ 4.5. Who tends to score high in buckpassing?
Figure 4. Examining the demographic distribution of buckpassing tendency: participants who have a high school education or below score higher in buckpassing (23% in high buckpassing level); participants who are aged under 18 and in their early 20s also tend to score higher in buckpassing. The y-axis is the proportion of participants in each buckpassing group.
Our analysis indicates that a buckpassing tendency significantly influences participants’ use of AI suggestions and reliance on AI. Thus we further explored the demographic distribution of buckpassing scores among our participants to gain insights into which demographic groups exhibit a higher buckpassing tendency.
Figure 4 shows the distribution of buckpassing levels among different education levels and age groups. Age is treated as a categorical variable in this descriptive analysis to obtain a more granular understanding of age group differences. Teenagers and young adults in our sample tended to score higher in buckpassing compared to participants who were aged above 35. Participants who obtain a lower level of education (high school education or below) also exhibited a higher buckpassing tendency compared to those with a higher level of education (pursuing or having obtained college or graduate education).
@@ 5. Discussion
Individuals’ decision-making patterns have been shown to influence their decision outcomes and quality of life (Kim et al., 2022; Deniz, 2006), underlining the importance of understanding how people vary in making decisions. However, despite the increasing role of AI in decision-making contexts, how individual cognitive decision-making processes influence the use of AI technology remains unknown. Our study offers evidence of this influence; we discuss our findings and consider their implications for designing AI systems and explainable AI research below.
@@
In this work, we extend the current understanding of individuals’ interactions with AI-generated information and factors behind these interactions in a setting where AI assistance is available but optional. Consistent with our hypotheses, we found that people who tend to defer decision-making patterns to others (score high in buckpassing) are more likely to seek out suggestions from an AI and report relying on it, though they spend less time reading AI’s explanations. While the difference can be described as a small to medium effect—those who score low on buckpassing viewed the AI’s decisions 25% of the time versus 33% for those who scored high on buckpassing—the effect of this difference can be profound when AI decision-making tools are frequently used and the performance of AI tools is not high. As we observed in our data, participants scored lower in accuracy when exposed to inaccurate AI decisions and higher when exposed to correct AI decisions. The difference may also gain importance with the number of decisions to be made as decision fatigue sets in (Chen et al., 2023) or as task difficulty or uncertainty increases (Vasconcelos et al., 2023; Salimzadeh et al., 2024), leading to a higher likelihood of overreliance.
@@
We encourage future work to examine these factors in various kinds of human-AI interactions. In addition, prior HCI research shows that individuals’ trust and reliance vary by their age, education levels, and perception of AI (Ma and Huo, 2023; Araujo et al., 2020). Our work extends this by showing that individuals’ buckpassing tendency has a marginal effect on AI use when accounting for these demographic covariates. While the significance of the effect decreases after controlling these factors, this is likely due to the correlation between these demographic factors and decision-making patterns (see Appendix Table 10 for the correlation matrix).
@@
What might explain these findings is the stress level and confidence people experience while making decisions. Mann et al. (1997)’s work has indicated that these decision-making tendencies can be seen as different kinds of coping mechanisms during decision conflict. Individuals with a strong buckpassing tendency often experience high psychological stress during decision-making tasks, while vigilant decision-makers experience a lower level of stress (Mann et al., 1997). Experiencing high psychological stress while making decisions might propel people to seek out an immediate relief by turning to AI suggestions. Mann et al. (1998) also identified that individuals who score high in buckpassing tend to have lower confidence in their own decision-making. This could potentially propel participants in our study to seek AI’s suggestions when they were not confident about making the decisions themselves.
Interestingly, we did not observe any effect of vigilance or hypervigilance on participants’ tendency to see AI suggestions. This might be due to the low variance of participants on vigilance ($SD=2.31$) and hypervigilance ($SD=2.61$) compared to buckpassing pattern ($3.03$). However, we observed that participants who scored high in vigilance spent significantly more time seeing AI explanations. This suggests that increased vigilance goes hand in hand with careful reading of any explanations people are provided, likely as a tool for evaluating whether the AI’s decision can be reasonably trusted.
@@
AI-assisted decision-making could yield both positive and negative outcomes for people with a tendency for buckpassing. On the one hand, the rise of AI systems could ease the burden of information seeking and processing for individuals to make effective decisions, if the performance of these systems is reliable. As evidenced by prior work, making decisions with the help of AI could lead to more accurate decisions when the AI has high advice accuracy and quality (Vodrahalli et al., 2021; Tschandl et al., 2020) (though there are some exceptions where this was found not to be the case (Vaccaro and Waldo, 2019; Jacobs et al., 2021)). AI assistance can also lower mental demand (Buçinca et al., 2021). With highly reliable AI systems, AI suggestions could help relieve the decision-making stress experienced by people with a buckpassing tendency. On the other hand, given the current risks of AI hallucinations, where AI generates unfaithful or nonfactual content (Bubeck et al., 2023), and AI overconfidence (Zhou et al., 2024), where AI outputs sound confident yet are inaccurate (Zhou et al., 2024), buckpassers are more susceptible to inaccurate information generated by AI, potentially resulting in misleading decisions and negative decision outcomes.
The benefits and risks of AI-assisted decision-making might be exacerbated for some people and groups, as our study and other research show that decision-making patterns vary across demographics. Consistent with prior findings (Filipe et al., 2020), our study suggests that buckpassing tendency is more prevalent among younger people (especially individuals under 18) and people with high school education or below. This warrants greater intervention to guide these populations in their usage of AI systems. Prior work from Mann et al. (1998) indicates that people from East Asian countries (e.g., Japan, Hong Kong) exhibit a higher buckpassing tendency compared to people in Western countries (e.g., USA, Australia, New Zealand). East Asians have also been found to have a lower decision-making self-esteem, which is generally in line with a preference for making decisions in a group (Mann et al., 1998). These demographic groups may seek out information from the AI more often, benefiting more from AI systems that have high accuracy but also facing greater risks of using inaccurate AI-generated information.